OpenStudy (anonymous):

What is (2a+b-c)^8. Or at least what is the sum of the coefficients?

OpenStudy (anonymous):

my guess is \(4^8\) but i could be wrong

OpenStudy (tkhunny):

Good guess, but not quite. a = b = c = 1 (2a + b - c) = 2 \(2^{8}\) This principle is nicely demonstrated with smaller exponents. \((2a + b - c)^{1} = 2a + b - c = 2^{1}\) \((2a + b - c)^{2} = 4a^{2}+4ab-4ac+b^{2}-2bc+c^{2} = 2^{2} = 4\) Some very tedious induction would demonstrate it definitively.

OpenStudy (anonymous):

So this means that the answer would be\[2^{8}\]?

OpenStudy (tkhunny):

Didn't I say that already? I also decided that the induction wouldn't be all that tedious.

OpenStudy (anonymous):

Just making sure. Sorry

OpenStudy (tkhunny):

No, no. Didn't you see my big smile when I said that? :-)

OpenStudy (anonymous):

Yes. :) It is simply late and I have been doing math for a while, and I only saw the bottom half of the post...